OpenStudy (anonymous):

Solve the given equation. cos θ = sqrt3/2

OpenStudy (anonymous):

I got \[\theta =\pi/6+2kpi,\theta= 2\pi/6 +2kpi\]

OpenStudy (savvy):

principle value is pi/6

OpenStudy (anonymous):

yes

OpenStudy (savvy):

general value is\[2k \pi \pm \pi/6\]

OpenStudy (anonymous):

so i was right?

OpenStudy (anonymous):

Well maybe for the second part.

OpenStudy (savvy):

to some extent..... you wrote just + it could be a '-' too.... and i didn't get how did that pi/6 became 2pi/6

OpenStudy (anonymous):

Wouldn't I just add 2pi/6 to pi/6 to get 2pi/6??